Add: 3mn, – 5mn, 8mn, – 4mn Add: 3mn, – 5mn, 8mn, – 4mn

Add: 
3mn, – 5mn, 8mn, – 4mn
Add: 
3mn, – 5mn, 8mn, – 4mn

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

When term have the same algebraic factors, they are like terms. Then, we have to add the like terms
= 3mn + (-5mn) + 8mn + (- 4mn)
= 3mn–5mn + 8mn–4mn
= mn (3–5 + 8 - 4)
= mn (11 - 9)
= mn (2)
= 2mn

Thanks